OpenStudy (anonymous):

Mark is graduating from college in six months, but he will need a loan in the amount of $7,038 for his last semester. He may either receive an unsubsidized Stafford Loan with an interest rate of 6.8%, compounded monthly, or his parents may obtain a PLUS Loan with an interest rate of 7.8%, compounded monthly. The Stafford Loan has a grace period of six months from the time of graduation. Which loan will have a higher balance at the time of repayment and by how much?

OpenStudy (anonymous):

7038(1+0.068/12)^12=7,531.79 7038(1+0.078/12)^6=7316.98 7531.79-7316.98=214.80 So the stafford loan has a higher balance by 214.80. This isnt in my answer choice it says the opposite.

OpenStudy (anonymous):

If an amount of $5,000 is deposited into a savings account at an annual interest rate of 5%, compounded monthly, the value of the investment after 10 years can be calculated as follows... P = 5000. r = 5/100 = 0.05 (decimal). n = 12. t = 10. If we plug those figures into the formula, we get: A = 5000 (1 + 0.05 / 12) ^ 12(10) = 8235.05. So, the investment balance after 10 years is $8,235.05.

OpenStudy (anonymous):

u can do it like this

OpenStudy (anonymous):

I did do it like that the aanswers given just do not match up